arnet is requested by his math teacher to make a pyramid whose base has a side which measure 12 cm. what would it be its volume required is 720 cubic centimeter
Answer by Cromlix(4381)

I am assuming this is a square based pyramid.
You are seeking the height that will give you
a volume of 720 cm^3
Volume of a square based pyramid = 1/3 Area of base x height
Your base being square has an area of 144 cm^2
Therefore:
Volume = 1/3 Area x height
720 cm^3 = 1/3 x 144 x height
Multiply both sides by 3
2160cm^3 = 144 x height
Divide both sides by 144
Height = 15 cm
Hope this helps.
:-)